• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Lindgren

    Log entry time 16:10:31 on June 9,2010

    Entry number 325418

    keyword=Shift Summary

    Day shift
    Richard Lindgren SL
    Nicholas Lubinsky TO
    Amber McGreary 3rd

    8:00 Hall A in Contolled Access
    RHRS Down
    Jack in the Hall
    Waiting for Run coordinator for evaluation

    Noticed target vacuum getting slightly worse with beam off
    Turned cooler off and heater on
    Seems to have stabilized the vacuum

    Sending 3rd worker to MCC to sign High Radiation Watch papers

    9:30 Amber McCreary on Rad Watch after long discussion,

    Rad watchers were assigned throughout the day while HRSR was worked on.
    At 13:00 jack segal said the HRSR was still down and not repaired yet.

    day shift ended